# Quantum Electrodynamics of graphene Landau levels in a deep-sub-wavelength hyperbolic phonon polariton cavity

### v1

@@ II.2 Theory of Landau levels in Graphene and their coupling to a subwavelength cavity
II.3 The polaritonic spectrum and the vacuum Rabi frequency
Alternatively, one can diagonalize the total Hamiltonian ^$\hat{\cal{H}}_{\rm tot}=\hat{\cal{H}}_{\rm LLs}+\hat{\cal{H}}_{\rm C}+\hat{\cal{H}}_{\rm cav}+\hat{\cal{H}}_{\rm int}$. In the single excitation manifold, electron-hole excitations between LLs can be treated as bosonic quasiparticles ciuti2005 . In this limit, the total Hamiltonian reduces to a quadratic bosonic Hopfield Hamiltonianhopfield1958 , which can be diagonalized exactly. The polariton energies obtained in this procedure are identical to those that one finds from the zeros of dynamical dielectric function $\epsilon_{\rm RPA}({q}_{\parallel},\omega)$.
Fig. 3 shows the electronic spectrum at the neutrality point $N_{\rm F}=0$. Panel (a) presents this spectrum as a function of the wavevector $q_{\parallel}$, while panel (b) displays the spectrum as a function of $(\hbar v_{\rm F}/\ell_{B})$. The latter spectrum can be experimentally obtained by varying the external magnetic field $B$ zhang2005 . Inside the Reststrahlen bands (defined by $\omega^{\rm T}_{i}<\omega<\omega^{\rm L}_{i}$, with these frequencies indicated by black dashed lines), the LL transitions hybridize with HPP cavity modes. The vacuum Rabi frequencies of the system represent the interaction strength between a selected LL transition and a cavity mode. As detailed in Section V, one of these couplings can reach up to $10\%$ of the cavity frequency, placing our system in the ultra-strong coupling regimeKockum19 ; geiser2012 . Furthermore, a single LL transition hybridizes with multiple cavity modes, a regime known as super-strong coupling kuzmin2019 . This region of the spectrum is mainly controlled by the LL transitions encoded in ${\chi}^{(0)}_{\rm nn}({q}_{\parallel},\omega)$ and by the quantum Green’s function associated with the cavity modes ${g}^{(0)}_{\rm q}({q}_{\parallel},0,0,\omega)$, corresponding to the expansion in Eq. (3). Outside the Reststrahlen bands, LL transitions experience a $q_{\parallel}$-dependent positive frequency shift, known as the depolarization shiftGiuliani_and_Vignale ; shtrichman2001 , leading to finite dispersion as apparent in Fig. 3(a). Panel (b) shows that, even off-resonance, the LL transitions do not coincide with the non-interacting case (indicated by white solid and dashed lines). This shift originates from the frequency-independent screening contribution of the Green’s function $g_{\infty}^{(0)}({q}_{\parallel},0,0)$ rather than from the HPP poles, as it occurs outside the Reststrahlen bands, where $\omega\neq\omega_{{q}_{\parallel},n_{z},s}$ This effect is largely due to the confinement induced by the two metallic mirrors. Indeed, this effect is significantly suppressed in the limit $L_{z}\to\infty$ (see Fig. 8(c) in Sect. V and the related discussion), which corresponds to removing the metallic mirrors and filling the entire space with hBN.
Fig. 4 shows the electronic spectrum for a doped system with $N_{\rm F}=5$, corresponding to a filling factor $\nu=22$, as $\nu=4N_{\rm F}+2$. At finite charge density $N_{\rm F}\neq 0$ and strong magnetic fields, a plasmonic excitation appears in the system, usually referred to as magnetoplasmon. Magnetoplasmons Giuliani_and_Vignale ; roldan2009 ; muravev2016 arise in doped semiconductors under a magnetic field due to the coupling between plasma oscillations of free carriers and their cyclotron resonance. In Fig. 4(a) we show the spectrum without the contribution of the quantum mechanical modes (which is obtained by setting $g_{\rm q}^{(0)}({q}_{\parallel},0,0,\omega)=0$), revealing the expected magnetoplasmon dispersionroldan2009 . Fig. 4(b) is obtained by taking into account the full cavity contribution. Here, the magnetoplasmon strongly hybridizes with the cavity modes achieving a Rabi spitting of $\approx 25\%$ of the cavity frequency. This significant coupling arises from the well-known fact that most of the oscillator strength, which quantifies the ability of a system to absorb or emit electromagnetic radiation, is associated with the magnetoplasmon resonanceGiuliani_and_Vignale . Finally, we note that in the $L_{z}\to\infty$ limit the Rabi splitting shown in Fig. 4(c) remains of a similar order. Although the coupling between LLs decreases with reduced compression, the corresponding increase in the density of cavity modes compensates, resulting in a splittingDai2015 ; Kumar2015 comparable to the case in which $L_{z}$ is finite. This effect is peculiar of subwavelength cavities, as transverse contributions from the vector potential $\hat{\bm{A}}$ vanish in the limit $L_{z}\to\infty$ ashida2023 . Similar considerations hold for the cavity spectrum, $S_{\rm cav}({q}_{\parallel},\omega)$. The details of this section are provided in Sec. V.
@@ III Details on “Theory of the Green’s function of HPP cavities”
In this Section, we outline the key steps and methods used to derive the scalar Green’s function for HPP cavities. A quantum framework is presented to relate the Green’s function to cavity field quantization, with explicit expressions connecting the scalar potential’s vacuum fluctuations to mode coefficients. Finally, the static and dynamic components of the Green’s function are distinguished, ensuring a comprehensive description of both resonant and off-resonant effects in the cavity’s electromagnetic response. These results form the theoretical foundation for the study of hPP cavities and their interaction with embedded materials.
@@
is generally finite. In the high-frequency limit, far from resonant poles, allowing us to consider $g_{\infty}^{(0)}(\bm{r},\bm{r}^{\prime})$ as a real function. However, the real and imaginary parts of the difference $g^{(0)}(\bm{r},\bm{r}^{\prime},\omega)-g_{\infty}^{(0)}(\bm{r},\bm{r}^{\prime})$, being causal response function that goes to zero for $\omega\to\infty$, satisfy the Kramers-Kronig relations Giuliani_and_Vignale :
@@ III.1 A quantization scheme for sub-wavelength cavities
using Eq. (42), we find:
@@
By applying Eq. (31) to the hBN cavity, we obtain Eq. (3) while the explicit expression of the normalization factors $N_{q_{\parallel},n_{z},s}$ is given in Eq. (LABEL:eq:N_q). The Green’s function can be expressed in terms of the coefficients $\phi_{q_{\parallel},n_{z},s}(z)$ as:
$\begin{split}&g_{\rm q}^{(0)}(q_{\parallel},z,z^{\prime},\omega)=\sum_{n_{z},s}\phi_{{q}_{\parallel},n_{z},s}(z)\phi_{{q}_{\parallel},n_{z},s}(z^{\prime})\times\\ &\times\left(\frac{1}{\hbar\omega-\hbar\omega_{q_{\parallel},n_{z},s}+\mathrm{i}\hbar\eta}-\frac{1}{\hbar\omega+\hbar\omega_{q_{\parallel},n_{z},s}+\mathrm{i}\hbar\eta}\right)~{}.\\ \end{split}$
Additionally, the vacuum fluctuations of the field are related to the coefficients $\phi_{q_{\parallel},n_{z},s}(z)$:
@@
Hence, the coefficients $\phi_{q_{\parallel},n_{z},s}(z)$ can be interpreted as a measure of the density of vacuum fluctuations. Moreover, these coefficients also appear in the Green’s function, as shown in Eq. (LABEL:eqn:green_function_hBN_expansion_quantum), indicating that the intensity of a scalar field generated by placing a test charge depends on the vacuum fluctuations of the field in the absence of any charge, represented by $\phi_{q_{\parallel},n_{z},s}(z)$. It thus is useful to express $\phi_{q_{\parallel},n_{z},s}(z)$ in the following form:
@@
IV Details on “Theory of Landau levels in Graphene and their coupling to a subwavelength cavity”
@@
In this Section, we review the LLs in graphene under a uniform magnetic field $B$ and their coupling to a subwavelength cavity. We derive the Hamiltonian governing LLs, emphasizing the eigenstates and associated energies $\varepsilon_{\lambda,n}$, which depend on the magnetic field strength via the magnetic length $\ell_{B}$. Subsequently, we use eigenstates and eigenenergies to calculate the density–density response function $\chi^{(0)}_{nn}(q_{\parallel},\omega)$. These findings will be useful later to calculate the polaritonic specturm.
@@ V Details on “The polaritonic spectrum and the vacuum Rabi frequency”
In this Section, we derive and diagonalize the Hopfield Hamiltonian to describe the polaritonic spectrum of our system. The Hamiltonian includes contributions from LLs, Coulomb interactions, HPP cavity modes, and light-matter coupling. By introducing bright modes, the LL Hamiltonian is expressed in terms of bosonic operators, and the interaction with cavity modes is described using Rabi frequencies, defined by vacuum fluctuation coefficients and by matrix elements of the density operator.
@@ V.1 Diagonalizing the Hopfield Hamiltonian
For convenience, we decompose this Hamiltonian into contributions from electrons and holes, such that ^$\hat{\mathcal{H}}_{\rm LLs}=\hat{\mathcal{H}}^{(+)}_{\rm LLs}+\hat{\mathcal{H}}^{(-)}_{\rm LLs}$ where
@@
Thus, the LLs Hamiltonian simplifies to:
^$\begin{split}\hat{\mathcal{H}}_{\rm LLs}\approx&\sum_{n^{\prime},\lambda^{\prime}\in\mathcal{N}_{\rm unocc}}\sum_{n,\lambda\in\mathcal{N}_{\rm occ}}\sum_{q_{x}}\big{(}\varepsilon_{\lambda,n}-\varepsilon_{\lambda^{\prime},n^{\prime}}\big{)}\\ &\times\hat{b}^{\dagger}_{n,\lambda,n^{\prime},\lambda^{\prime}}(q_{x})\hat{b}_{n,\lambda,n^{\prime},\lambda^{\prime}}(q_{x})~{},\end{split}$ (75)
@@ V.2 The vacuum Rabi frequencies
[…]
Figure 7: (Color online) Panels (a,d) display the ratio $R_{q_{\parallel},n_{z},s,-,0,+,1}=\Omega_{q_{\parallel},n_{z},s,-,0,+,1}/\omega_{q_{\parallel},n_{z},s}$, where $\Omega_{q_{\parallel},n_{z},s,-,0,+,1}$ denotes the Rabi frequency for the Landau level (LL) transition $(0,+1)$, and $\omega_{q_{\parallel},n_{z},s}$ is the cavity mode frequency. This ratio is shown as a function of $q_{\parallel}$ (in ${\rm nm}^{-1}$) for different values of $n_{z}$, with $L_{z}=30{\rm nm}$ and $B=10{\rm T}$. Panels (b,e) display the same ratio as a function of $L_{z}$ (in ${\rm nm}$) for different $n_{z}$ for a fixed $\bar{q}_{\parallel}=0.1{\rm nm^{-1}}$ and $B=10{\rm T}$. Panels (c,f) present the Rabi frequency as a function of $\hbar v_{\rm F}/\ell_{B}$ (in ${\rm eV}$) for different $n_{z}$. Panels (a,b,c) show these quantities for the upper Reststrahlen band using blue dotted curves, while panels (d,e,f) present the lower Reststrahlen band with red dashed curves. The line thickness diminishes as the quantum number $n_{z}$ increases.
@@
Eq. (87) can be interpreted as describing the response of a system composed of harmonic modes with frequencies $\omega_{q_{\parallel},n_{z},s}$, which are coupled to other harmonic modes with frequencies $(\varepsilon_{\lambda^{\prime},n^{\prime}}-\varepsilon_{\lambda,n})/\hbar$. The coupling strength between these modes is characterized by the Rabi frequency $\hbar\Omega_{q_{\parallel},{n}z,{s},{\lambda},{n},{\lambda}^{\prime},{n}^{\prime}}$.
@@
In standard C-QED dalibard1982 ; DDB_arxiv_2023 ; ciuti2005 , it is well known that the Rabi frequency is proportional to the product of the vacuum fluctuations of the electric field $\sqrt{\langle\hat{E}^{2}\rangle}$ and the dipole moment of the atom in the cavity, $d$, such that $\Omega_{\rm R}\sim d\sqrt{\langle\hat{E}^{2}\rangle}$. Eq. (88) generalizes this concept to non-local ($q_{\parallel}\neq 0$), sub-wavelength cavities interacting with a 2D material. As seen from Eq. (88), the Rabi frequency is proportional to the product of the vacuum fluctuation coefficients $\phi_{{q}_{\parallel},n_{z},s}(0)$ and the density matrix elements $\sqrt{{\mathcal{M}_{n^{\prime},\lambda^{\prime},n,\lambda}({q}_{\parallel})}}$. Therefore, to maximize the light-matter coupling, one can separately optimize the vacuum fluctuations and the density matrix elements. Specifically, $\phi_{{q}_{\parallel},n_{z},s}(0)$ reaches its maximum when $q_{\parallel}L_{z}\sim 1$, while $\sqrt{{\mathcal{M}_{n^{\prime},\lambda^{\prime},n,\lambda}({q}_{\parallel})}}$ peaks when $q_{\parallel}\ell_{B}\sim 1$. This leads to the conclusion that for optimal coupling, all the relevant length scales must be comparable, such that $q_{\parallel}\sim 1/\ell_{B}\sim 1/L_{z}$.
@@
In Fig. 7, we present the ratio $R_{q_{\parallel},n_{z},s,-,0,+,1}\equiv(\Omega_{q_{\parallel},n_{z},s,-,0,+,1}/\omega_{q_{\parallel},n_{z},s})$, where $\Omega_{q_{\parallel},n_{z},s,-,0,+,1}$ is the Rabi frequency associated with the LL transition $(0,+1)$ and $\omega_{q_{\parallel},n_{z},s}$ is the cavity frequency as a function of different parameters. In Fig. 7(a,d), the ratio $R_{q_{\parallel},n_{z},s,-,0,+,1}$ is plotted as a function of $q_{\parallel}$ (in ${\rm nm}^{-1}$) for various $n_{z}$. This plot reveals a non-monotonic behavior in $q_{\parallel}$, where each mode exhibits a specific $q_{\parallel}$ that maximizes the Rabi frequency. Similarly, Fig. 7(b,e) displays this quantity as a function of $L_{z}$ (in ${\rm nm}$) for different $n_{z}$. This also shows a non-monotonic dependence on $L_{z}$, which contrasts with the naive expectation that the Rabi frequency would scale as $1/\sqrt{L_{z}}$. This scaling is valid only for $q_{\parallel}L_{z}\gg 1$, whereas for small $L_{z}$, the coupling vanishes. Finally, Fig. 7(c,f) shows the Rabi frequency as a function of $\hbar v_{\rm F}/\ell_{B}$, which can be controlled experimentally by adjusting the external magnetic field. At low values, specifically around $\hbar v_{\rm F}/\ell_{B}\approx 0.01~{}{\rm eV}$, increasing the magnetic field—and thus this ratio—leads to a stronger coupling. However, the coupling saturates when $\hbar v_{\rm F}/\ell_{B}$ reaches approximately $0.15~{}{\rm eV}$ or higher.
@@
In Fig. 7(a-f), we observe that the maximum ratio $R_{q_{\parallel},n_{z},s,-,0,+,1}$, achieved for $n_{z}=0$, exceeds $10\%$ for the upper Reststrahlen band and approaches $10\%$ for the lower band. This ratio places the system firmly within the ultra-strong coupling regime, where the coupling strength is a significant fraction of the cavity frequency Kockum19 . Additionally, a single LL hybridizes with multiple cavity modes, as the Rabi frequencies of many different modes remain non-negligible. This situation characterizes a regime known as super-strong coupling kuzmin2019 .

### v2

@@ II.3 The polaritonic spectrum and the vacuum Rabi frequency
II.4 Relation to prior works
As discussed in the Introduction, many theoretical works on cavity quantum materials rely either on a phenomenological uniform vector potential or on the vector potential derived for Fabry-Pérot cavities, which host modes that fulfill the diffraction limit. These approaches are often applied to sub-wavelength cavities using a phenomenological compression factor $A$ gao2020 ; schlawin2019 ; wang2019 ; nguyen2023 ; zezyulin2022 ; lenk2022 ; arwas2023 ; pavosevic2022 ; Dag2024 or by fitting the vacuum field to experimental values.
Ref. andolina2024, demonstrated that, in the generalized Coulomb gauge, sub-wavelength cavities — characterized by a large compression factor $A=(V/\lambda^{3}_{\rm d})\ll 1$ — exhibit an electric field with a predominantly electrostatic nature, described by the scalar potential $\hat{\phi}$, while the magnetic field remains negligible. Consequently, in such cavities, the effective electron-electron interactions, obtained by tracing out the cavity degrees of freedom, are primarily density-density interactions governed by the dielectric tensor $\epsilon_{i,j}(\bm{r},\omega)$. In contrast, the transverse current-current interactions, which dominate in Fabry-Pérot cavities, are not significantly enhanced in sub-wavelength cavities, as they originate from Ampère’s law and thus depend only on the magnetic permeability $\mu(\bm{r},\omega)$ rather than on the dielectric tensor. However, in sub-wavelength cavities, the magnetic permeability remains approximately that of free space, $\mu(\bm{r},\omega)\approx 1$, preventing a substantial enhancement of these interactions. This example underscores the necessity of developing a new theoretical framework capable of microscopically describing the vacuum electromagnetic field in sub-wavelength cavities and its coupling with matter.
Here, we developed this framework by quantizing sub-wavelength cavities. For any generic sub-wavelength cavity, our theory enables the quantization of the electric field and thus the calculation of the mode amplitude coefficients, $\phi_{{q}_{\parallel},n_{z},s}(z)$, given in Eq. (3), which represent the vacuum fluctuations of the electric field, as well as the vacuum Rabi frequencies $\Omega_{q_{\parallel},{n}_{z},{s},{\lambda},{n},{\lambda}^{\prime},{n}^{\prime}}$ in Eq. (8). Additionally, our calculations reveal the renormalization of the electrostatic Coulomb interaction, represented by $g^{(0)}_{\infty}({q}_{\parallel},0,0)$. We applied our theory to electrons in graphene subject to a strong perpendicular magnetic field $B$ and coupled to the sub-wavelength modes of an HPP cavity. Ref. ashida2023, investigated graphene in a magnetic field embedded in a cavity where the mirrors are made of hBN. Despite the apparent similarity, that study focused solely on transverse cavity modes, considering only the excitations of the vector potential $\hat{\bm{A}}$, while neglecting the longitudinal cavity modes $\hat{\phi}$ and the Coulomb interaction $g^{(0)}_{\infty}({q}_{\parallel},0,0)$. However, as demonstrated in Ref. andolina2024, , when the field is confined below the diffraction limit, $V\ll\lambda_{\rm d}^{3}$, the electrostatic contribution encoded in $g^{(0)}({q}_{\parallel},0,0,\omega)$ are dominant. Thus, our analysis is valid in the regime of strong confinement and highly non-local wavevectors considered in this article, whereas the theory in Ref. ashida2023, applies to the complementary limit of low wavevectors, where retardation effects due to the finite speed of light become crucial.
In the future, it would be interesting to apply our theory to investigate a 2DEG in a sub-wavelength split-ring cavity, motivated by pioneering experiments on the quantum Hall effect in such systems appugliese2022 ; enkner2024 ; enkner2024b . Indeed, most of previous theoretical works ciuti2021 ; rokaj2022 ; rokaj2023 that have considered this setup employed a single uniform transverse mode, modeled as $\hat{\bm{A}}={{A}}_{0}\bm{\epsilon}\left(\hat{a}+\hat{a}^{\dagger}\right)$, where $\hat{a},\hat{a}^{\dagger}$ are the photonic operators, $\bm{\epsilon}$ is the polarization and ${A}_{0}=\sqrt{2\pi\hbar c^{2}/V_{\rm eff}\omega_{c}}$ is the vacuum cavity field in a Fabry-Pérot cavity, $\omega_{c}$ is the cavity frequency and $V_{\rm eff}=A\lambda^{3}_{\rm d}$ is the effective cavity volume rescaled by the use of the phenomenological compression factor $A$. While this approach may capture the main phenomenology, it neglects the non-local structure of the cavity mode enkner2024b , lacks a microscopic prediction of the vacuum field without resorting to the phenomenological compression factor or fitting experimental values, and omits frequency-independent electrostatic effects, which in our theory are encoded in $g_{\infty}(\bm{r},\bm{r}^{\prime})$. We believe that a fully microscopic treatment of the vacuum field in sub-wavelength cavities, applying our theory to a 2DEG in a sub-wavelength split-ring cavity while accounting for both the non-local field $\hat{\phi}(\bm{r})$ and the renormalized Coulomb potential $g_{\infty}(\bm{r},\bm{r}^{\prime})$, is therefore necessary to accurately describe these experiments.
@@ III Details on “Theory of the Green’s function of HPP cavities”
In this Section, we outline the key steps and methods used to derive the scalar Green’s function for HPP cavities. Furthermore, we develop a quantization scheme for the electric field in a sub-wavelength cavity by employing the cavity scalar Green’s function. This approach enables us to relate the vacuum fluctuations of the scalar potential to the amplitude coefficients of each mode. Finally, we decompose the scalar Green’s function into its static and dynamical components, allowing us to distinguish between resonant effects and off-resonant contributions associated with screening.
@@
is generally finite. In the high-frequency limit, far from resonant poles, the function $g_{\infty}^{(0)}(\bm{r},\bm{r}^{\prime})$ can be assumed to be a real function. However, the real and imaginary parts of the difference $g^{(0)}(\bm{r},\bm{r}^{\prime},\omega)-g_{\infty}^{(0)}(\bm{r},\bm{r}^{\prime})$, being causal response function that goes to zero for $\omega\to\infty$, satisfy the Kramers-Kronig relations Giuliani_and_Vignale :
@@ III.1 A quantization scheme for sub-wavelength cavities
By means of Eq. (45) and Eqs. (III,46) we find the following decomposition:
@@
By applying Eq. (34) to the hBN cavity, we obtain Eq. (3) while the explicit expression of the normalization factors $N_{q_{\parallel},n_{z},s}$ is given in Eq. (LABEL:eq:N_q). The Green’s function can be expressed in terms of the amplitude coefficients $\phi_{q_{\parallel},n_{z},s}(z)$ as:
$\begin{split}&g_{\rm q}^{(0)}(q_{\parallel},z,z^{\prime},\omega)=\sum_{n_{z},s}\phi_{{q}_{\parallel},n_{z},s}(z)\phi_{{q}_{\parallel},n_{z},s}(z^{\prime})\times\\ &\times\left(\frac{1}{\hbar\omega-\hbar\omega_{q_{\parallel},n_{z},s}+\mathrm{i}\hbar\eta}-\frac{1}{\hbar\omega+\hbar\omega_{q_{\parallel},n_{z},s}+\mathrm{i}\hbar\eta}\right)~{}.\\ \end{split}$ (50)
Additionally, the vacuum fluctuations of the field are related to the amplitude coefficients $\phi_{q_{\parallel},n_{z},s}(z)$:
@@
Hence, the amplitude coefficients $\phi_{q_{\parallel},n_{z},s}(z)$ can be interpreted as a measure of the density of vacuum fluctuations. Moreover, these coefficients also appear in the Green’s function, as shown in Eq. (LABEL:eqn:green_function_hBN_expansion_quantum), indicating that the intensity of a scalar field generated by placing a test charge depends on the vacuum fluctuations of the field in the absence of any charge, represented by $\phi_{q_{\parallel},n_{z},s}(z)$. It thus is useful to express $\phi_{q_{\parallel},n_{z},s}(z)$ in the following form:
@@
IV Details on “Theory of Landau levels in Graphene and their coupling to a sub-wavelength cavity”
@@
In this Section, we review the LLs in graphene under a uniform magnetic field $B$ and their coupling to a sub-wavelength cavity. We derive the Hamiltonian governing LLs, emphasizing the eigenstates and associated energies $\varepsilon_{\lambda,n}$, which depend on the magnetic field strength via the magnetic length $\ell_{B}$. Subsequently, we use eigenstates and eigenenergies to calculate the density–density response function $\chi^{(0)}_{nn}(q_{\parallel},\omega)$. These findings will be useful later to calculate the polaritonic specturm.
@@ V Details on “The polaritonic spectrum and the vacuum Rabi frequency”
In this Section, we derive and diagonalize the Hopfield Hamiltonian to describe the polaritonic spectrum of our system. By introducing bright modes, the LLs Hamiltonian is expressed in terms of bosonic operators and coupled to the cavity Hamiltonian via the interaction Hamiltonian. Once the Hopfield Hamiltonian is derived, the vacuum Rabi frequencies can be identified as the coupling constant associated with light-matter interaction. This approach establishes a direct connection between the vacuum Rabi frequencies, the vacuum fluctuation coefficients of the cavity, and the matrix elements of the electronic density operator. Additionally, we discuss the depolarization shift, linking it to the non-resonant contribution of the scalar Green’s function. Finally, we identify a regime where only a single cavity mode is resonant with a single LL transition, allowing for an effective description within a simplified two-transition model.
@@ V.1 Diagonalizing the Hopfield Hamiltonian
We now want to describe a system with many electrons occupying the Fermi Sea. For convenience, we decompose this Hamiltonian into contributions from electrons and holes, such that ^$\hat{\mathcal{H}}_{\rm LLs}=\hat{\mathcal{H}}^{(+)}_{\rm LLs}+\hat{\mathcal{H}}^{(-)}_{\rm LLs}$ where
@@
Thus, we can write an effective Hamiltonian for the electron-hole excitations above the Fermi sea as:
^$\begin{split}\hat{\mathcal{H}}^{\rm eff}_{\rm LLs}&=\sum_{n^{\prime},\lambda^{\prime}\in\mathcal{N}_{\rm unocc}}\sum_{n,\lambda\in\mathcal{N}_{\rm occ}}\sum_{q_{x}}\big{(}\varepsilon_{\lambda^{\prime},n^{\prime}}-\varepsilon_{\lambda,n}\big{)}\\ &\times\hat{b}^{\dagger}_{n,\lambda,n^{\prime},\lambda^{\prime}}(q_{x})\hat{b}_{n,\lambda,n^{\prime},\lambda^{\prime}}(q_{x})~{},\end{split}$ (79)
@@ V.2 The vacuum Rabi frequencies
Figure 8: (Color online) Panels (a,c) display the ratio $R_{q_{\parallel},n_{z},s,-,0,+,1}=\Omega_{q_{\parallel},n_{z},s,-,0,+,1}/\omega_{q_{\parallel},n_{z},s}$, where $\Omega_{q_{\parallel},n_{z},s,-,0,+,1}$ denotes the Rabi frequency for the Landau level (LL) transition $(0,+1)$, and $\omega_{q_{\parallel},n_{z},s}$ is the cavity mode frequency. This ratio is shown as a function of $q_{\parallel}$ (in ${\rm nm}^{-1}$) for different values of $n_{z}$, with $L_{z}=30{\rm nm}$ and $B=10{\rm T}$. Panels (b,d) present the Rabi frequency as a function of $\hbar v_{\rm F}/\ell_{B}$ (in ${\rm eV}$) for different $n_{z}$. Panels (a,b) show these quantities for the upper Reststrahlen band using blue dotted curves, while panels (c,d) present the lower Reststrahlen band with red dashed curves. The line thickness diminishes as the quantum number $n_{z}$ increases.
@@
Eq. (91) can be interpreted as describing the response of a system composed of harmonic modes with frequencies $\omega_{q_{\parallel},n_{z},s}$, which are coupled to other harmonic modes with frequencies $(\varepsilon_{\lambda^{\prime},n^{\prime}}-\varepsilon_{\lambda,n})/\hbar$. The coupling strength between these modes is characterized by the Rabi frequency $\hbar\Omega_{q_{\parallel},{n}z,{s},{\lambda},{n},{\lambda}^{\prime},{n}^{\prime}}$. Eq. (92) demonstrates how the vacuum Rabi frequency is determined by the fluctuations of both the electronic density and the electric field, extending the standard concept of cavity quantum electrodynamics (C-QED) dalibard1982 ; DDB_arxiv_2023 ; ciuti2005 to non-local ($q_{\parallel}\neq 0$) sub-wavelength cavities interacting with a 2D material. Indeed, as evident from Eq. (92), the Rabi frequency is proportional to the product of the vacuum fluctuation coefficients $\phi_{{q}_{\parallel},n_{z},s}(0)$ and the density matrix elements $\sqrt{{\mathcal{M}_{n^{\prime},\lambda^{\prime},n,\lambda}({q}_{\parallel})}}$. Therefore, to maximize the light-matter coupling, one can separately optimize the vacuum fluctuations and the density matrix elements. Specifically, $\phi_{{q}_{\parallel},n_{z},s}(0)$ reaches its maximum when $q_{\parallel}L_{z}\sim 1$, while $\sqrt{{\mathcal{M}_{n^{\prime},\lambda^{\prime},n,\lambda}({q}_{\parallel})}}$ peaks when $q_{\parallel}\ell_{B}\sim 1$. This leads to the conclusion that for optimal coupling, all the relevant length scales must be comparable, such that $q_{\parallel}\sim 1/\ell_{B}\sim 1/L_{z}$.
@@
In Fig. 8, we present the ratio $R_{q_{\parallel},n_{z},s,-,0,+,1}\equiv(\Omega_{q_{\parallel},n_{z},s,-,0,+,1}/\omega_{q_{\parallel},n_{z},s})$, where $\Omega_{q_{\parallel},n_{z},s,-,0,+,1}$ is the Rabi frequency associated with the LL transition $(0,+1)$ and $\omega_{q_{\parallel},n_{z},s}$ is the cavity frequency as a function of different parameters. In Fig. 8(a,c), the ratio $R_{q_{\parallel},n_{z},s,-,0,+,1}$ is plotted as a function of $q_{\parallel}$ (in ${\rm nm}^{-1}$) for various $n_{z}$. This plot reveals a non-monotonic behavior in $q_{\parallel}$, where each mode exhibits a specific $q_{\parallel}$ that maximizes the Rabi frequency. Fig. 8(b,d) shows the Rabi frequency as a function of $\hbar v_{\rm F}/\ell_{B}$, which can be controlled experimentally by adjusting the external magnetic field. At low values, specifically around $\hbar v_{\rm F}/\ell_{B}\approx 0.01~{}{\rm eV}$, increasing the magnetic field—and thus this ratio—leads to a stronger coupling. However, the coupling saturates when $\hbar v_{\rm F}/\ell_{B}$ reaches approximately $0.15~{}{\rm eV}$ or higher.
@@
In Fig. 4 (a-b) and Fig. 8(a-d), we observe that the maximum ratio $R_{q_{\parallel},n_{z},s,-,0,+,1}$, achieved for $n_{z}=0$, exceeds $10\%$ for the upper Reststrahlen band and approaches $10\%$ for the lower band, placing our system in the ultra-strong coupling regime, where the coupling strength is a significant fraction of the cavity frequency Kockum19 .